OpenStudy (anonymous):

what is the circumfrence of 12mm diameter

OpenStudy (31356):

Welcome to Open Study! I think you went to the wrong section, this is the Music section. You will get better assistance in the Math area. Good luck! :D

OpenStudy (praxer):

radius = 6mm, hence circumference is 2pir =2*pi*r =2*3.14*6